OpenStudy (anonymous):

how do i find the inverse of f(x)=9/11x

OpenStudy (aum):

\[ f(x) = \frac{9}{11}x ~~ \text{ OR } ~~ f(x) = \frac{9}{11x} \]

OpenStudy (aum):

Which one of the above is the correct problem?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

h

OpenStudy (aum):

y = 9 / (11x) solve for x: multiply both sides by 1x y * x = 9/11 divide both sides by y: x = 9 / (11y) interchange x and y: y = 9/(11x) Inverse = 9/(11x)

OpenStudy (anonymous):

i dont get it...

OpenStudy (aum):

If you are given a function f(x) and asked to find the inverse, these are the steps: 1. Replace f(x) by y. Here we rewrote f(x)=9/(11x) as y = 9/(11x) 2. Solve for x. Here, x = 9/(11y) 3. Interchange x and y. y = 9/(11x) 4. Replace y with \(f^{-1}(x)\). Therefore the inverse function is: \(\large f^{-1}(x) = \frac{9}{11x}\).

OpenStudy (aum):

Here the inverse function is same as the original function. This happens whenever the function is symmetrical about the y = x line.

OpenStudy (anonymous):

alright, its just that i have to write the question like this:

OpenStudy (aum):

???

OpenStudy (anonymous):

f^-1(x) = ax +b/ cx + d)

OpenStudy (aum):

\[ \large f^{-1}(x) = \frac{9}{11x} = \frac{0x + 9}{11x + 0} \]a = 0, b = 9, c = 11, d = 0